OpenStudy (anonymous):

Find the volume of a cone with a diameter of 7 inches and a height that is 4 times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (5 points) Select one: a. 29.31 in3 b. 87.92 in3 c. 117.23 in3 d. 179.50 in3

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

use v=1/3*(pi)*r^2*h

OpenStudy (anonymous):

what would the radius be in this one

OpenStudy (sidsiddhartha):

radius=3.5

OpenStudy (anonymous):

v=1/3*(3.14)*(3.5)^2*(?)

OpenStudy (anonymous):

i dont know what the height would be

OpenStudy (sidsiddhartha):

h=4*r=4*3.5

OpenStudy (anonymous):

so the height is 4

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

Find the volume of a cone with a diameter of 7 inches and a height that is 4 times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (5 points) Given; diameter of the cone d= 7 in hence radius r = 7/2 = 3.5 in Since height is 4times the radius therefore, height h= 4*r = 4*3.5 = 14 in Volume of cone is given as: \[\huge V = \frac{1}{3} \pi r^2 h= \frac{1}{3} \times 3.14 \times (3.5)^2 \times 14\] \[\huge \rightarrow V = \frac{1}{3} \times 3.14 \times 12.25 \times 14\] \[\huge \rightarrow V = \frac{1}{3} \times 38.465 \times 14\] \[\huge \rightarrow V = \frac{1}{3} \times 538.51\] \[\huge \rightarrow V = 179.50333 \approx 179.50 \] Hence the volume of the given cone is approximately equal to 179.50 cubic inches. Thus option d. 179.50 in3 is the correct answer. @prowrestler